OpenStudy (anonymous):

A cylinder has a radius of 3m and total surface area of 42pi find the height

OpenStudy (anonymous):

\[2\pi r ^{2} + 2\pi rh = 42 \pi\] \[18\pi + 6\pi h = 42 \pi\] \[6\pi h = 24\pi\] h=4 m